# Approximation and bounding techniques for the Fisher-Rao distances

### v2

@@ 6 Tight lower bounds on Fisher-Rao distances from isometric embeddings
Although there are many ways to design a variety of upper bounds on the Fisher-Rao distance, it is more difficult to design lower bounds. The classic technique for designing a lower bound on the Fisher-Rao distance of a statistical model $\mathcal{M}$ is find an isometric embedding of $\mathcal{M}$ as a submanifold of a higher-dimensional manifold for which the Riemannian distance is known in closed form. We first illustrate this principle for the Fisher-Rao distance on the categorical distributions in §6.1 (multinomial distributions with a single trial also called multinoulli distributions by analogy to the link of Bernoulli distributions with Binomial distributions), and then discuss the technique for submodels of the family of multivariate normal distributions.
@@ 6.2 Case of totally geodesic submanifolds
Property 3 (Comparing Fisher-Rao distances under isometric embededings)
@@ 7.4 Calvo and Oller’s Fisher-Rao isometric embeddings onto the SPD matrix cone
Property 4 (Calvo-Oller lower bound)
@@ 9 Conclusion: Summary and some open problems
When the Fisher-Rao metric $g$ is Hessian we get a canonical pair of dual torsion-free flat affine connections $\nabla$ and $\nabla^{*}$ which mid-connection +*$\frac{\nabla+\nabla^{*}}{2}$ coincides with the Levi-Civita metric connection induced by the Fisher metric. We can check whether the Fisher information matrix given in any parameterization is Hessian or not using the test of Proposition 1. In dimension $1$ and $2$, all analytic metrics are Hessian metrics [7]. We noticed that the dually flat connections $\nabla$ and $\nabla^{*}$ associated to a Hessian metric may not be necessarily Amari’s expected $\alpha$-connections as this can be attested for some univariate elliptical models [72] like the family of Cauchy distributions for which all Amari’s expected $\alpha$-connections coincide with the non-flat Fisher-Rao connection. Yet the Fisher Cauchy metric is Hessian and admits two canonical flat connections [84]. To implement Algorithm 3 for Fisher Hessian metrics, we reported a canonical upper bound on their Fisher-Rao distances (Proposition 8), and show how to design lower bounds from isometric embeddings in §6. In particular, we considered Calvo and Oller isometric embeddings [27] for multivariate elliptical distribution families and discussed how to implement the various approximation and upper bounds for those families in §7.5. We also proposed a fast alternative metric distance called the Birkhoff-Calvo-Oller distance for multivariate elliptical distribution families (Definition 1).
@@
Problem 5.
Given a Hessian metric $g(\theta)$ for =$\mathcal{M}=\{p_{\theta}{\ :\ }\theta\in\Theta\}$, report a parameterization $\xi$ and a Legendre-type convex function $F(\xi)$ such that $G(\xi)=\nabla^{2}F(\xi)$, i.e., $g_{ij}(\xi)=\frac{\partial}{\partial\xi_{i}}\frac{\partial}{\partial\xi_{j}}\,F(\xi)$.

### v3

@@ 6 Tight lower bounds on Fisher-Rao distances from isometric embeddings
Although there are many ways to design upper bounds on the Fisher-Rao distance, it seems more difficult to design lower bounds. The classic technique for designing a lower bound on the Fisher-Rao distance of a statistical model $\mathcal{M}$ is find an isometric embedding of $\mathcal{M}$ as a submanifold of a higher-dimensional manifold for which the Riemannian distance is known in closed form.
We first illustrate this principle for the Fisher-Rao distance on the categorical distributions in §6.1 (multinomial distributions with a single trial also called multinoulli distributions by analogy to the link of Bernoulli distributions with Binomial distributions), and then discuss the technique for submodels of the family of multivariate normal distributions.
@@ 6.1 Isometric embedding of the Fisher-Rao multinoulli manifold
Using Nash embedding theorem [92, 58, 56], a $m$-dimensional Riemannian manifold $(M,g)$ can always be embedded as a submanifold of the Euclidean manifold $(E,g_{E})$ of dimension $n\leq 2m$. Let $f:M\rightarrow E$ be the embedding function. Then we have
$\rho_{g}(p_{1},p_{2})\geq\rho_{E}(f(p_{1}),f(p_{2}))=\|f(p_{1})-f(p_{2})\|_{2}.$
Let us notice that computing explicitly the embedding function $f$ is a difficult problem in general although some algorithmic works tackle this problem [120]. Hilbert [59] proved that the hyperbolic upper plane manifold (Poincaré hyperbolic metric for $m=2$) cannot be fully embedded in the $n=3$-dimensional Euclidean space. Only a partial band of the upper plane can be isometrically embedded onto the pseudo-sphere [102] of $\mathbb{R}^{3}$.
@@ 6.2 Case of totally geodesic submanifolds
Property 4 (Comparing Fisher-Rao distances under isometric embededings)
@@ 7.4 Calvo and Oller’s Fisher-Rao isometric embeddings onto the SPD matrix cone
Property 5 (Calvo-Oller lower bound)
@@ 9 Conclusion: Summary and some open problems
When the Fisher-Rao metric $g$ is Hessian we get a canonical pair of dual torsion-free flat affine connections $\nabla$ and $\nabla^{*}$ which mid-connection $\frac{\nabla+\nabla^{*}}{2}$ coincides with the Levi-Civita metric connection induced by the Fisher metric. In dimension $1$ and $2$, all analytic metrics are Hessian metrics [8]. We noticed that the dually flat connections $\nabla$ and $\nabla^{*}$ associated to a Hessian metric may not be necessarily Amari’s expected $\alpha$-connections as this can be attested for some univariate elliptical models [80] like the family of Cauchy distributions for which all Amari’s expected $\alpha$-connections coincide with the non-flat Fisher-Rao connection. Yet the Fisher Cauchy metric is Hessian and admits two canonical flat connections [94]. To implement Algorithm 3 for Fisher Hessian metrics, we reported a canonical upper bound on their Fisher-Rao distances (Proposition 8), and show how to design lower bounds from isometric embeddings in §6. In particular, we considered Calvo and Oller isometric embeddings [28] for multivariate elliptical distribution families and discussed how to implement the various approximation and upper bounds for those families in §7.5. We also proposed a fast alternative metric distance called the Birkhoff-Calvo-Oller distance for multivariate elliptical distribution families (Definition 1).
@@
Given a Riemannian metric $g$ expressed using parameterization $\lambda$ (i.e., given $g(\lambda)$ on a single global chart), how to find, if possible, a parameterization $\theta$ and a potential convex function $F(\theta)$ such that $g(\theta)=\nabla^{2}F(\theta)$ (i.e., check whether the geometric metric $g$ is Hessian or not).
Problem 4.
@@
Appendix A Symbolic computing of Fisher information
We use the computer algebra software Maxima (https://maxima.sourceforge.io/) for computing symbolically the Fisher information matrix. For example, the code snippet below calculates the Fisher information of bivariate normal distributions [106]:
kill(all)\$
v:matrix([x],[y]);
mu:matrix([a],[b]);
Sigma: matrix([c,d],[d,e]);
assume(c>0);assume(e>0);
assume(determinant(Sigma)>0);
/* pdf of a bivariate normal distribution */
pdf:1/(2*%pi*sqrt(determinant(Sigma)))*exp(-(1/2)*transpose(v-mu).invert(Sigma).(v-mu));
/* Fisher information matrix */
hessian(-log(pdf),[a,b,c,d,e]);
integrate(integrate(pdf*%,x,minf,inf),y,minf,inf);
We can check that the expectation of the score vanishes:
/* check that the expectation of the score is zero */
derivative(log(pdf),a,1);
integrate(integrate(pdf*%,x,minf,inf),y,minf,inf);
The Fisher information matrix for some cases of trivariate normal distributions was reported in [50].